The number of sigma and pi-bonds in 1-butene-yne are:
(a) 5 sigma and 5 pi
(b) 7 sigma and 3 pi
(c) 8 sigma and 2dpi
(d) 5 sigma and 4 pi

Grade:upto college level

3 Answers

Navjyot Kalra
9 years ago
Hello Student,
Please find the answer to your question
(b). CH2 = CH – C ≡ CH;
No of σ bonds = 2 + 1 + 1 + 1 / 1 + 1 = 7;
No of π bonds = 1 + 2 = 3

Thanks
Navjyot Kalra
askIITians Faculty
Akash Rao
6 years ago
Please find the answer to your question(b). CH2 = CH – C ≡ CH;No of σ bonds = 2 + 1 + 1 + 1 / 1 + 1 = 7;No of π bonds = 1 + 2 = 3
ankit singh
3 years ago
CH2 = CH – C ≡ CH;
No of σ bonds = 2 + 1 + 1 + 1 / 1 + 1 = 7;
No of π bonds = 1 + 2 = 3
thanks and regards
ankit singh
 manit bhopal
askiiitians forum answering faculty